OpenStudy (javeneno):

I will medal and fan!!!! The product of 3 and a number is 12 more than the sum of the number and 26. What is the number?

OpenStudy (javeneno):

Plz help

OpenStudy (javeneno):

@saifoo.khan

OpenStudy (javeneno):

i will medal and fan

OpenStudy (javeneno):

@Hero

OpenStudy (javeneno):

@Directrix

OpenStudy (javeneno):

@campbell_st

OpenStudy (javeneno):

hi

OpenStudy (javeneno):

i understand the expression as 3x = 12 + (26y)

OpenStudy (campbell_st):

let the number be n \[3 \times n = (n + 26) + 12\] now simplify then solve this for n

OpenStudy (javeneno):

i still dont quit get it

OpenStudy (javeneno):

would it be the same as 3 x n = n + 38 @campbell_st

OpenStudy (javeneno):

@Abhisar

OpenStudy (campbell_st):

that's correct... now solve it 3n = n + 38

OpenStudy (javeneno):

i understand thanks @campbell_st